# Dust destruction by the supernova remnant forward shock in a turbulent interstellar medium

### v1

@@ 3.1 Low density ISM
Figure 5: $16\times 16$ pc zoom-in on a filament of the $n=1\,\mathrm{cm^{-3}}$ and $\mathcal{M}=1$ simulation at $t=0\,\mathrm{yr}$ and $t=10\,\mathrm{kyr}$. The filament cannot withstand the shock efficiently but is destroyed leaving behind finger like structures that arise from the forward shock front into the SNR interior. Furthermore, some high density cloudlets are formed, as circled in the image. The arrow indicates the approximate direction of the forward shock.
Looking closer, a few high density cloudlet structures can be seen that are a few pixels wide and have higher densities than the initial filaments ($n>100\,\mathrm{cm^{-3}}$, e.g. the ones circled in Fig. 5). These cloudlets result from compressed material of the processed filament and can shield some dust from destruction as they take longer to be disrupted than the lower density surroundings.
Although the filaments do not withstand the energetic forward shock but get processed to lower density fingers and eventually destroyed, they are able to shield some dust in comparison to the homogeneous case in the most destructive, first 10 kyr because it takes longer for them to be penetrated and processed by the shock than in the homogeneous case. Roughly 14% of silicates and 8% of carbonaceous dust grains are shielded by the filaments in the supersonic turbulent Mach number case, decreasing with turbulence strength to 4% of silicate and 3% of carbonaceous dust for a subsonic Mach number as can be seen in Fig. 6. This supports the conclusions of previous papers with a longer evolution time that find that filaments or inhomogeneities in general are able to shield dust from destruction(Hu2019; Kirchschlager2024a; Dedikov2025). However, the destroyed dust mass remains on the same order of magnitude for all Mach number simulations since the forward shock encounters more dust in the high Mach number cases.
@@
This happens due to the turbulent ISM structures in the high Mach number cases that lead to an irregular expansion of the SNR. While it can rapidly move through the voids that have a very low density, the forward shock is slowed down more in higher density regions. Since filaments are spatially thin in comparison to the voids, the shock can evolve through lower density areas around them (c.f. Korolev2015; Slavin2017; Dedikov2025). Thus, after 10 kyr, the forward shock undergoes greater expansion in the high Mach number ISM than in the homogeneous case and is also more efficiently slowed down on average across the entire shock, due to its low velocity in the filaments.
@@
The velocity of the forward shock, which can be estimated by averaging the gas velocity of the shock cells traced by a momentum threshold, is therefore still 388 km s-1 in the homogeneous case, whereas it slowed down to 251 km s-1 in the supersonic turbulence case after 10 kyr, as given in Table 1. Since kinetic sputtering is directly dependent on the forward shock velocity, a comparison of the simulations at a similar end shock velocity instead of similar end time would also be a valuable comparison. This was done in Appendix F, where all simulations are compared at different evolution times but similar average shock velocities of 387–390 km s-1. For the low density simulations, the spread in dust destruction efficiency between the different Mach numbers increases to 26% for the silicate grains and 16% for the carbonaceous grains, however, the high Mach number simulations show significantly less destroyed dust masses in this comparison, reflecting the trend that the filaments can shield the dust from destruction.
@@
Regardless of the comparison criterion of the simulations (Table 1 or 4) and the studied parameter set, silicates are always at least 12% more efficiently destroyed than carbonaceous grains in a low density environment. This difference can generally be explained by the higher sputtering yields of silicates that stem from their higher bulk density [^3] The bulk density of silicate and carbonaceous grains is 3.3 g cm-3 and 2.2 g cm-3, respectively (Tielens1994; Jones1994).. Thereby, the silicate grains are also less efficiently accelerated by the forward shock, resulting in higher relative dust-to-gas velocities and thus more efficient kinetic sputtering at the shock front. However, other mechanisms are also affected. Due to their higher bulk density, silicates are less numerous than carbonaceous grains since we consider the same dust sizes and total dust mass for both species, resulting in less grain-grain collisions and a lower sputtering efficiency due to a lower total grain surface area. Furthermore, silicates have lower relative grain-to-grain velocities when they are accelerated by the forward shock so that fragmentation and vaporization in general are less efficient for silicates than for carbonaceous grains. Moreover, the size-dependency of sputtering is considered in Paperboats (see Kirchschlager2019), which arises since smaller grain sizes are used in comparison to the experimental sputtering yields which considered a semi-infinite target (Jurac1998; SerraDiaz-Cano2008). This correction was updated in our study to consider hydrogen ions of the ISM to impinge on dust grains as projectiles instead of oxygen ions as used previously by Kirchschlager2022; Kirchschlager2024a. It generally leads to less silicate grains but more small carbonaceous grains being destroyed, supporting a better survival of silicates. For the low density simulations, however, the generally higher sputtering yields and more efficient kinetic sputtering during the acceleration of the grains seem to be more important than the higher sputtering efficiency due to a higher total grain surface area, the corrections of the size-dependent sputtering, and the higher fragmentation and vaporization efficiencies.
@@
In total, more than 1.3 M of silicate dust is destroyed in each low density simulation in the first 10 kyr, which is higher than the maximum amount of dust ($\sim$1 M) produced in SNRs as inferred from observations. This threshold is, however, not reached in the case of carbonaceous grains yet, of which slightly less than 1 M is destroyed in the first 10 kyr. Although this seems to imply that SNRs are net dust destroyers for an ISM with predominantly silicate grains but can be net producers for a carbonaceous dust-dominated environment, the results of our simulations are only lower limits of the dust destruction, since we only investigate the first 10 kyr of SNR evolution. At 10 kyr, the dust destruction efficiency of all low density simulations has not yet plateaued, meaning that the dust is still being efficiently destroyed. Furthermore, the forward shock is still fast and hot at 10 kyr (¿240 km s-1) so that it is reasonable to expect efficient dust destruction by kinetic and thermal sputtering beyond 10 kyr. If the so far neglected reverse shock dust destruction would also be taken into account, it is probable that SNRs destroy more dust than they produce in this low density environment, even if the ISM is dominated by carbonaceous dust.
@@ 3.2 Intermediate density ISM
The SNR forward shock in the intermediate density environment ($n=10\,\mathrm{cm^{-3}}$, central column of Fig. 3) is slowed down more efficiently than in the low density environment, ending at 198–226 km s-1 (Table 1) and only reaching sizes of 16–28 pc. In total, 58–92% or 4.4–5.6 M of silicate dust, and 46–77% or 3.5–4.7 M of carbonaceous dust is destroyed. Thus, carbonaceous grains are again more robust than silicate grains with similar differences of at least 12% more destruction of silicates as in the low density simulations.
@@
With more than 3.5 M of destroyed dust for every intermediate density simulation, significantly more dust is being destroyed than in the low density cases. The destruction efficiency is also higher since collisions of the hot and high-velocity ions with the ISM dust grains occur more often, resulting in more efficient thermal and kinetic sputtering. Grain-grain collisions are still not very efficient in the homogeneous and subsonic Mach number cases, leading to differences of $\sim$1%. They make a bigger impact of 2.9–12% for the more turbulent cases of this ISM density as they are typically more efficient in higher density media such as filaments. For all cases, similar to the low density environment, the grain-size distribution with and without grain-grain collisions is similar for grains larger than 5 nm but is altered for smaller grains (see Appendix E). In general, the final grain-size distribution is altered in a similar way as in the low density cases, having more smaller grains destroyed than larger grains. However, as grain-grain collisions become effective in the high Mach number simulations, more and more large grains are being destroyed.
@@
Compared to the low density ISM, the filaments in the intermediate density transonic and supersonic cases have a higher number density, mostly on the order of 1000 cm-3. Thus, they can better withstand the forward shock that itself is generally more decelerated due to the higher average ISM density. When the filaments are encountered by the forward shock, they are not immediately destroyed, but partly swept up and dragged along with the shock, so that a thin and dense shell-like structure can form, which can cool down more efficiently (see Fig. 3). However, most of the filaments are still eventually destroyed, leaving behind finger-like structures that expand into the ejecta, similar to Fig. 5, and high density cloudlets.
@@
Because of the longer filament survival times and the lower penetration velocity, the dust is shielded much better in the dense regions of the higher Mach number cases, independent of the dust species, as shown in Fig. 6. Thus, we can observe a large difference in dust destruction efficiencies between homogeneous and strongly turbulent ISM of up to 34% for the silicate grain simulations and 31% for the carbonaceous grain simulations, decreasing with Mach number to 0.9% dust destruction difference with silicates and 0.6% difference with carbonaceous grains. As a result, also less dust mass is being destroyed by the SNR in the higher Mach number simulations, although it expands to larger sizes.
@@
At similar end shock velocities of 387–390 km s-1 instead of similar end times, this trend remains visible and accounts for up to 39% and 34% of dust destruction difference for the silicate and carbonaceous grains, respectively (Appendix F). The amount of destroyed dust mass also has a larger spread between the Mach number cases and remains higher than in the low density ISM, however, the dust destruction difference between intermediate and low density ISM cases decreases slightly.
@@ 3.3 High density ISM
In contrast to the intermediate and low density cases, the forward shock in the high density case is being compressed to a thin and dense shell from $\sim$3–4 kyr onward. This happens because larger amounts of gas are swept up than in the lower density cases, the shock is being decelerated more to $v<130\,\mathrm{km\,s^{-1}}$, and the SNR interior cools down enough. The formation of the shell can be seen well in the high density, homogeneous ISM simulation in Fig. 3, in which the SNR interior is empty whereas the shell has a very high density and starts to exhibit Vishniac thin-shell modes (Vishniac1983; Vishniac1994), which are inhibited by the irregular expansion of the forward shock in the higher Mach turbulence cases.
@@
From the time the dense shell forms onward, grain-grain collisions become very efficient. They account for a maximum difference of 36% in dust destruction at 10 kyr, which is significantly more than in our low and intermediate density models. Due to the high density of the shell, the forward shock starts to cool down more efficiently, and thus slows down more so that the effectiveness of both thermal and kinetic sputtering decreases. At a certain critical velocity, which is dependent on the grain composition, the kinetic sputtering yield drops steeply, making it even less efficient (Tielens1994; Nozawa2006). Hence, after 3–4 kyr, the sputtering-only simulations start to have a shallow slope in dust destruction (Fig. 4). The simulations with all destruction mechanisms included continue to have high destruction efficiencies. Thus, with the onset of the shell, the significance of the destruction mechanisms alters since sputtering becomes inefficient whereas grain shattering gains in effectiveness.
@@
Due to this change, the destruction efficiencies of carbonaceous and silicate grains in the high density ISM are only 3% different at most. For the weakly turbulent and homogeneous cases, carbonaceous grains are even slightly more efficiently destroyed than silicates. This partly results from the higher shattering efficiency of carbonaceous grains in comparison to silicate grains, as discussed in Sect. 3.1, and becomes evident from the larger differences of at least 6% between the carbonaceous and silicate grain simulations with only sputtering considered, shown in Table 1. However, as described in Sect. 3.1, several mechanisms play a role in destroying more carbonaceous than silicate grains in these particular simulations.
@@
The grain-size distribution also changes more significantly in these high density cases because larger dust grains are more efficiently destroyed by shattering than smaller grains. The processed grain-size distribution is therefore shifted to smaller grain sizes compared to the initial MRN distribution, as discussed in Appendix E.
@@
Since sputtering becomes significantly less efficient with the formation of the shell, the trend of a higher dust destruction efficiency with higher ISM density also discontinues with the $n=100\,\mathrm{cm^{-3}}$ cases although more dust mass is being destroyed because more dust is encountered. While the forward shock dust destruction efficiency in a high density and low Mach number ISM is still higher than in the intermediate density medium at 3 kyr, it is roughly 4–5% lower than in the intermediate density cases at 10 kyr. If only sputtering is considered, almost all other Mach number cases also decrease in destruction efficiency in the high density ISM compared to the intermediate density cases after 10 kyr, by up to 21%.
@@
However, a comparison at the same end shock velocities as in Appendix F still shows the trend of more dust destruction with a higher average ISM density because the shell has not yet formed during the modeled period. Thus, a higher dust fraction and more dust mass is destroyed in the high density ISM than in the intermediate density cases. We note that this comparison has to be judged carefully since the SNR shock does not encounter many filamentary structures in the higher Mach number simulations yet, and thus, mostly evolves in a void with a density $n\lesssim$10 cm-3, rather resembling an intermediate or low density case. Thus, the dust destruction in the higher Mach number simulations is also much lower than in the homogeneous ISM with large differences of up to 36%.
@@
Smaller differences are found after 10 kyr of evolution time. The destruction efficiency from the homogeneous to the strongly turbulent ISM case is only reduced by 11% for the silicate grains and 14% for the carbonaceous grains, as seen in Fig. 6. This is less than in the intermediate density cases and similar to the low density cases although the filaments are able to slow down the shock more efficiently and almost stop it at high densities. The forward shock shape at the end of our simulations is similar to the initial filament structures, as shown in Fig. 7, and almost no finger-like shapes are formed but some dust-shielding cloudlets occur. However, due to the shell formation in the high density cases and the subsequent shift of significant destruction mechanisms, the shielding ability of the filaments decreases. Since shattering is more efficient in higher density media, it shrinks the difference in dust destruction between the homogeneous and strongly turbulent cases by destroying dust in dense filaments more efficiently than in the homogeneous and weakly turbulent runs. Thus, at roughly 4 kyr, the dust destruction efficiency difference between weak and strong turbulence is still on the same order of the intermediate density cases with over 30%, but decreases again after the shell forms.
@@
Figure 7: $9\times 9$ pc zoom-in on a filament of the $n=100\,\mathrm{cm^{-3}}$ and $\mathcal{M}=3$ simulation at $t=0$ yr and $t=10$ kyr. The dense filament almost stops the shock and dictates where it will evolve. The initial filament shape looks nearly identical to the forward shock shape at the end. The arrow indicates the approximate direction of the forward shock.
@@
In total, SNRs always destroy more than 9 M of dust in their first 10 kyr of evolution in a high density ISM, which is a magnitude more than what SNe have been observed to produce at most, making them significant net dust destroyers in this environment. An even higher dust destruction would be found if the considered ejecta cooling would be less efficient than our rather strong oxygen cooling so that the SNR would expand more and the shell would form later on, resulting in more dust being encountered and thus destroyed and a longer period of efficient sputtering.

### v2

@@ 3.1 Low-density ISM
Figure 5: $16\times 16$ pc zoomed-in image of a filament of the $n=1\,\mathrm{cm^{-3}}$ and $\mathcal{M}=1$ simulation at $t=0\,\mathrm{yr}$ and $t=10\,\mathrm{kyr}$. The filament cannot withstand the shock efficiently, but is destroyed leaving behind finger-like structures that arise from the forward shock front into the SNR interior. Furthermore, some high-density cloudlets are formed (circled in black). The arrow indicates the approximate direction of the forward shock.
A few high-density cloudlet structures can be seen that are a few pixels wide and have higher densities than the initial filaments ($n>100\,\mathrm{cm^{-3}}$, e.g., the ones circled in Fig. 5). These cloudlets result from compressed material of the processed filament and can shield some dust from destruction as they take longer to be disrupted than the lower density surroundings.
Although the filaments do not withstand the energetic forward shock but get processed to lower density fingers and eventually destroyed, they are able to shield some dust in comparison to the homogeneous case in the most destructive, first 10 kyr because it takes longer for them to be penetrated and processed by the shock than in the homogeneous case. Roughly 14% of silicates and 8% of carbonaceous dust grains are shielded by the filaments in the supersonic turbulent Mach number case, decreasing with turbulence strength to 4% of silicate and 3% of carbonaceous dust for a subsonic Mach number as can be seen in Fig. 6. This supports the conclusions of previous papers with a longer evolution time that find that filaments or inhomogeneities in general are able to shield dust from destruction(Hu et al. 2019; Kirchschlager et al. 2024a; Dedikov and Vasiliev 2025). However, the destroyed dust mass remains on the same order of magnitude for all Mach number simulations since the forward shock encounters more dust in the high Mach number cases.
@@
This happens due to the turbulent ISM structures in the high Mach number cases that lead to an irregular expansion of the SNR. While it can rapidly move through the voids that have a very low density, the forward shock is slowed down more in higher density regions. Since filaments are spatially thin in comparison to the voids, the shock can evolve through lower density areas around them (cf. Korolev et al. 2015; Slavin et al. 2017; Dedikov and Vasiliev 2025). Thus, after 10 kyr, the forward shock undergoes greater expansion in the high Mach number ISM than in the homogeneous case and is also more efficiently slowed down on average across the entire shock, due to its low velocity in the filaments.
@@
The velocity of the forward shock, which can be estimated by averaging the gas velocity of the shock cells traced by a momentum threshold, is therefore still 388 km s-1 in the homogeneous case, whereas it slows down to 251 km s-1 in the supersonic turbulence case after 10 kyr, as given in Table 1. Since kinetic sputtering is directly dependent on the forward shock velocity, a comparison of the simulations at a similar end shock velocity instead of similar end time would also be a valuable comparison. This was done in Appendix F, where all simulations are compared at different evolution times but similar average shock velocities of 387–390 km s-1. For the low-density simulations, the spread in dust destruction efficiency between the different Mach numbers increases to 26% for the silicate grains and 16% for the carbonaceous grains; however, the high Mach number simulations show significantly less destroyed dust mass in this comparison, reflecting the trend that the filaments can shield the dust from destruction.
@@
Regardless of the comparison criterion of the simulations (Table 1 or 4) and the studied parameter set, silicates are always at least 12% more efficiently destroyed than carbonaceous grains in a low-density environment. This difference can generally be explained by the higher sputtering yields of silicates that stem from their higher bulk density. [^3] The bulk density of silicate and carbonaceous grains is 3.3 g cm-3 and 2.2 g cm-3, respectively (Tielens et al. 1994; Jones et al. 1994). Thus, the silicate grains are also less efficiently accelerated by the forward shock, resulting in higher relative dust-to-gas velocities and thus more efficient kinetic sputtering at the shock front. However, other mechanisms are also affected. Due to their higher bulk density, silicates are less numerous than carbonaceous grains since we consider the same dust sizes and total dust mass for both species, resulting in fewer grain-grain collisions and a lower sputtering efficiency due to a lower total grain surface area. Furthermore, silicates have lower relative grain-to-grain velocities when they are accelerated by the forward shock so that fragmentation and vaporization in general are less efficient for silicates than for carbonaceous grains. Moreover, the size-dependency of sputtering is considered in Paperboats (see Kirchschlager et al. 2019), which arises since smaller grain sizes are used in comparison to the experimental sputtering yields which considered a semi-infinite target (Jurac et al. 1998; Serra Díaz-Cano and Jones 2008). This correction was updated in our study to consider hydrogen ions of the ISM to impinge on dust grains as projectiles instead of oxygen ions as used previously by Kirchschlager et al. (2022, 2024a). It generally leads to fewer silicate grains but more small carbonaceous grains being destroyed, supporting a better survival of silicates. For the low-density simulations, however, the generally higher sputtering yields and more efficient kinetic sputtering during the acceleration of the grains seem to be more important than the higher sputtering efficiency due to a higher total grain surface area, the corrections of the size-dependent sputtering, and the higher fragmentation and vaporization efficiencies.
@@
In total, more than 1.3 M of silicate dust is destroyed in each low-density simulation in the first 10 kyr, which is higher than the maximum amount of dust ($\sim$1 M) produced in SNRs as inferred from observations. However, this threshold is not reached in the case of carbonaceous grains yet: slightly less than 1 M is destroyed in the first 10 kyr. Although this seems to imply that SNRs are net dust destroyers for an ISM with predominantly silicate grains but can be net producers for a carbonaceous dust-dominated environment, the results of our simulations are only lower limits of the dust destruction, since we only investigate the first 10 kyr of SNR evolution. At 10 kyr, the dust destruction efficiency of all low-density simulations has not yet plateaued, meaning that the dust is still efficiently destroyed. Furthermore, the forward shock is still fast and hot at 10 kyr (¿240 km s-1) so that it is reasonable to expect efficient dust destruction by kinetic and thermal sputtering beyond 10 kyr. If the so far neglected reverse shock dust destruction would also be taken into account, it is probable that SNRs destroy more dust than they produce in this low-density environment, even if the ISM is dominated by carbonaceous dust.
@@ 3.2 Intermediate-density ISM
The SNR forward shock in the intermediate-density environment ($n=10\,\mathrm{cm^{-3}}$, central column of Fig. 3) is slowed down more efficiently than in the low-density environment, ending at 198–226 km s-1 (Table 1) and only reaching sizes of 16–28 pc. In total, 58–92% or 4.4–5.6 M of silicate dust, and 46–77% or 3.5–4.7 M of carbonaceous dust is destroyed. Thus, carbonaceous grains are again more robust than silicate grains with similar differences of at least 12% more destruction of silicates as in the low-density simulations.
@@
With more than 3.5 M of destroyed dust for every intermediate-density simulation, significantly more dust is destroyed than in the low-density cases. The destruction efficiency is also higher since collisions of the hot and high-velocity ions with the ISM dust grains occur more often, resulting in more efficient thermal and kinetic sputtering. Grain-grain collisions are still not very efficient in the homogeneous and subsonic Mach number cases, leading to differences of $\sim$1%. They make a bigger impact of 2.9–12% for the more turbulent cases of this ISM density as they are typically more efficient in higher density media such as filaments. For all cases, similar to the low-density environment, the grain-size distribution with and without grain-grain collisions is similar for grains larger than 5 nm but is altered for smaller grains (see Appendix E). In general, the final grain-size distribution is altered in a similar way as in the low-density cases, having more smaller grains destroyed than larger grains. However, as grain-grain collisions become effective in the high Mach number simulations, more and more large grains are destroyed.
@@
Compared to the low-density ISM, the filaments in the intermediate-density transonic and supersonic cases have a higher number density, mostly on the order of 1000 cm-3. Thus, they can better withstand the forward shock that itself is generally more decelerated due to the higher average ISM density. When the filaments are encountered by the forward shock, they are not immediately destroyed, but partly swept up and dragged along with the shock, so that a thin and dense shell-like structure can form, which can cool down more efficiently (see Fig. 3). However, most of the filaments are still eventually destroyed, leaving behind finger-like structures that expand into the ejecta, similar to Fig. 5, and high-density cloudlets.
@@
Because of the longer filament survival times and the lower penetration velocity, the dust is shielded much better in the dense regions of the higher Mach number cases, independent of the dust species, as shown in Fig. 6. Thus, we can observe a large difference in dust destruction efficiencies between homogeneous and strongly turbulent ISM of up to 34% for the silicate grain simulations and 31% for the carbonaceous grain simulations, decreasing with Mach number to 0.9% dust destruction difference with silicates and 0.6% difference with carbonaceous grains. As a result, also less dust mass is destroyed by the SNR in the higher Mach number simulations, although it expands to larger sizes.
@@
At similar end shock velocities of 387–390 km s-1 instead of similar end times, this trend remains visible and accounts for up to 39% and 34% of dust destruction difference for the silicate and carbonaceous grains, respectively (Appendix F). The amount of destroyed dust mass also has a larger spread between the Mach number cases, and remains higher than in the low-density ISM; however, the dust destruction difference between intermediate- and low-density ISM cases decreases slightly.
@@ 3.3 High-density ISM
In contrast to the intermediate- and low-density cases, the forward shock in the high-density case is compressed to a thin and dense shell from $\sim$3–4 kyr onward. This happens because larger amounts of gas are swept up than in the lower density cases, the shock is decelerated more to $v<130\,\mathrm{km\,s^{-1}}$, and the SNR interior cools down enough. The formation of the shell can be seen well in the high-density, homogeneous ISM simulation in Fig. 3, in which the SNR interior is empty whereas the shell has a very high density and starts to exhibit Vishniac thin-shell modes (Vishniac 1983, 1994), which are inhibited by the irregular expansion of the forward shock in the higher Mach turbulence cases.
@@
From the time the dense shell forms onward, grain-grain collisions become very efficient. They account for a maximum difference of 36% in dust destruction at 10 kyr, which is significantly more than in our low- and intermediate-density models. Due to the high density of the shell, the forward shock starts to cool down more efficiently, and thus slows down more so that the effectiveness of both thermal and kinetic sputtering decreases. At a certain critical velocity, which is dependent on the grain composition, the kinetic sputtering yield drops steeply, making it even less efficient (Tielens et al. 1994; Nozawa et al. 2006). Hence, after 3–4 kyr, the sputtering-only simulations start to have a shallow slope in dust destruction (Fig. 4). The simulations with all destruction mechanisms included continue to have high destruction efficiencies. Thus, with the onset of the shell, the significance of the destruction mechanisms alters since sputtering becomes inefficient whereas grain shattering gains in effectiveness.
@@
Due to this change, the destruction efficiencies of carbonaceous and silicate grains in the high-density ISM are only 3% different at most. For the weakly turbulent and homogeneous cases, carbonaceous grains are even slightly more efficiently destroyed than silicates. This partly results from the higher shattering efficiency of carbonaceous grains in comparison to silicate grains, as discussed in Sect. 3.1, and becomes evident from the larger differences of at least 6% between the carbonaceous and silicate grain simulations with only sputtering considered, shown in Table 1. However, as described in Sect. 3.1, several mechanisms play a role in destroying more carbonaceous than silicate grains in these particular simulations.
@@
The grain-size distribution also changes more significantly in these high-density cases because larger dust grains are more efficiently destroyed by shattering than smaller grains. The processed grain-size distribution is therefore shifted to smaller grain sizes compared to the initial MRN distribution, as discussed in Appendix E.
@@
Since sputtering becomes significantly less efficient with the formation of the shell, the trend of a higher dust destruction efficiency with higher ISM density also discontinues with the $n=100\,\mathrm{cm^{-3}}$ cases although more dust mass is destroyed because more dust is encountered. While the forward shock dust destruction efficiency in a high-density and low Mach number ISM is still higher than in the intermediate-density medium at 3 kyr, it is roughly 4–5% lower than in the intermediate-density cases at 10 kyr. If only sputtering is considered, almost all other Mach number cases also decrease in destruction efficiency in the high-density ISM compared to the intermediate-density cases after 10 kyr, by up to 21%.
@@
However, a comparison at the same end shock velocities as in Appendix F still shows the trend of more dust destruction with a higher average ISM density because the shell has not yet formed during the modeled period. Thus, a higher dust fraction and more dust mass is destroyed in the high-density ISM than in the intermediate-density cases. We note that this comparison has to be judged carefully since the SNR shock does not encounter many filamentary structures in the higher Mach number simulations yet, and thus, mostly evolves in a void with a density $n\lesssim$10 cm-3, rather resembling an intermediate- or low-density case. Thus, the dust destruction in the higher Mach number simulations is also much lower than in the homogeneous ISM with large differences of up to 36%.
@@
Smaller differences are found after 10 kyr of evolution time. The destruction efficiency from the homogeneous to the strongly turbulent ISM case is only reduced by 11% for the silicate grains and 14% for the carbonaceous grains, as seen in Fig. 6. This is less than in the intermediate-density cases and similar to the low-density cases although the filaments are able to slow down the shock more efficiently and almost stop it at high densities. The forward shock shape at the end of our simulations is similar to the initial filament structures, as shown in Fig. 7, and almost no finger-like shapes are formed but some dust-shielding cloudlets occur. However, due to the shell formation in the high-density cases and the subsequent shift of significant destruction mechanisms, the shielding ability of the filaments decreases. Since shattering is more efficient in higher density media, it shrinks the difference in dust destruction between the homogeneous and strongly turbulent cases by destroying dust in dense filaments more efficiently than in the homogeneous and weakly turbulent runs. Thus, at roughly 4 kyr, the dust destruction efficiency difference between weak and strong turbulence is still on the same order of the intermediate-density cases with over 30%, but decreases again after the shell forms.
@@
Figure 7: $9\times 9$ pc zoomed-in image of a filament of the $n=100\,\mathrm{cm^{-3}}$ and $\mathcal{M}=3$ simulation at $t=0$ yr and $t=10$ kyr. The dense filament almost stops the shock and dictates where it will evolve. The initial filament shape looks nearly identical to the forward shock shape at the end. The arrow indicates the approximate direction of the forward shock.
@@
In total, SNRs always destroy more than 9 M of dust in their first 10 kyr of evolution in a high-density ISM, which is a magnitude more than what SNe have been observed to produce at most, making them significant net dust destroyers in this environment. An even higher dust destruction would be found if the considered ejecta cooling would be less efficient than our rather strong oxygen cooling so that the SNR would expand more and the shell would form later on, resulting in more dust being encountered and thus destroyed and a longer period of efficient sputtering.